The resistance of gold is 2.8 × 10^{-3} ? at a temperature of 20°C. A gold wire, 1.8 mm in diameter and 14 cm long, carries a current of 600 mA. The power dissipated in the wire is closest to:

A) 33 mW

B) 1.0 mW

C) 1.7 mW

D) 1.8 mW

E) 55 mW
Answer #1

Power dissipited in the wire

P=I2R

Given I=600mA=0.6A

R=2.8 × 10^{-3} Ω

P=0.62*2.8*10-3=1*10-3W= 1mW